the middle term expansion of (x^2+1/x^2+2)^n

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Parul,

(x^2 +1/x^2 +2)^n=(x+(1/x))2n

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Parul !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

Ashwin Muralidharan IIT Madras
12 years ago

Hi Parul,

 

x2+1/x2+2 = (x+1/x)2

 

So the given expression is (x+1/x)2n.

And the middle term would be 2nCnxn(1/x)n = 2nCn.

 

Hope this helps.

 

Best Regards,

Ashwin (IIT Madras).